A skydiver has a mass of 72 kg and a weight of 706 N. He is descending on a parachute and his speed is still increasing at 1.2 m/s2.

What is the magnitude and direction of the net force on the skydiver? Explain.

What is the magnitude and direction of the force that the parachute harness exerts on the skydiver? Explain.
Answer #1

There are two forces acting on the skydiver: 1) weight (w) in downward direction and 2) air resistance due to parachute in the upward direction

Since there is an acceleration of 1.2 m/s2 in downward direction. net force is in downward direction.

Fnet = ma = 72*1.2 = 86.4 N

Fnet = w - Fpara

Fpara = w - Fnet = 706 - 86.4 = 619.6 N

parachute harness exerts force in the upward direction on the skydiver